OpenStudy (anonymous):

Let \displaystyle f(x) = \frac{x-6}{x+2}. Find the open intervals on which f is increasing (decreasing). Then determine the x-coordinates of all relative maxima (minima). 1. f is increasing on the intervals 2. f is decreasing on the intervals 3. The relative maxima of f occur at x = 4. The relative minima of f occur at x =

OpenStudy (anonymous):

I know I have to get the derivative which is f'(x) = 8/(x+2)^2

OpenStudy (anonymous):

\[f(x) = \frac{x-6}{x+2}\]

OpenStudy (anonymous):

forget calc, just treat this rational function like you did in precalc class vertical asymptote \(x=-2\) horizontal asymptote \(y=1\) and picture should be easy

OpenStudy (anonymous):

assuming your derivative is right (i didn't check it) then since the numerator is 8 and the denominator is a square, that means the derivative is always positive (except at -2) so the function is always increasing

OpenStudy (anonymous):

you remember graphing stuff like this before you took calc?  it is still the same

OpenStudy (anonymous):

okay so what would I say for the increasing interval, cause it is not (-inf,inf) or (-inf,-2)U(-2,inf). So what would I say? And I do remember those graphs.

OpenStudy (anonymous):

i would say it is increasing everywhere except at -2 where it is undefined

OpenStudy (anonymous):

if you want interval notation \[(-\infty, -2)\cup (-2,\infty)\] which is a long winded way of saying "everything but -2"

OpenStudy (anonymous):

there is clearly no local max or min the derivative is never zero, the numerator of the derivative is a number

OpenStudy (anonymous):

Okay, thanks, I got it now.

OpenStudy (anonymous):

yw